OpenStudy (anonymous):

Find the coordinates of the midpoint of a segment with the endpoints (16,5) & (28, -13).

OpenStudy (owlcoffee):

You can find the mid-point of any segment given the end points of it with the following expression that represents a mid point "M" for any segment: \[M(\frac{ x_2+x_1 }{ 2 },\frac{ y_2+y_1 }{ 2 })\] It's just the sum of the coordinates of the end-points and then you divide them by two. So, fo the segments with the end-points (16,5) and (28,-13), the mid point "M" will have the following values: \[M(\frac{ 16+28 }{ 2 },\frac{ 5-13 }{ 2 })\] I'll leave the duty of calculating those fractions to you.